# Neural Collapse-Inspired Multi-Label Federated Learning under Label-Distribution Skew

### v2

@@ Introduction
Table 4: Comparison of related FL methods with ours on the ChestX-ray14 dataset (Wang et al.,, 2017).
Ablation Study
To evaluate the effectiveness of each component in the proposed method, we conducted an ablation study on the multi-label DermaMNIST dataset, with the client data distribution configured using $\beta=0.1$ and $\gamma=0.71$. The experimental results are presented in Table 5. As shown in the table, incorporating LADM for class-specific feature extraction, together with the predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, yields improvements of 0.43% in class-wise AUC and 7.34% in class-wise F1 score. Additional performance gains are achieved by introducing regularisation terms that enhance the model’s discriminative ability, resulting in further improvements of 3.31% in class-wise AUC and 3.43% in class-wise F1 score.
@@
query type query init macro-AUC macro-F1 AUC F1
@@
We also investigated the class-wise query mechanism within the attention-based LADM module. As reported in Table 6, fixed, well-designed class-wise queries consistently outperform learnable queries across most evaluation metrics. This outcome is intuitive in the context of label-skewed FL, where each client has a unique local data distribution. If the queries and classifier are learnable, the model may overfit to these local distributions, which can significantly deviate from the global distribution. Such overfitting introduces inconsistencies among client models and ultimately degrades the performance of the aggregated global model. To maintain consistent learning across clients with non-iid data, it is therefore preferable to use predefined, well-separated, fixed embeddings to guide the learning process rather than relying on learnable representations.
@@
To further analyse model behaviour, we present t-SNE visualisations of feature distributions from the test data in the latent feature space, generated under different model architectures and training strategies on the multi-label DermaMNIST dataset. The client data distribution is configured using $\beta=0.1$ and $\gamma=0.71$. As shown in Figure 3, when a conventional, learnable fully connected (FC) classifier is used, although the model achieves 83.95% class-wise AUC and 40.69% F1 score, the resulting feature representations exhibit poor clustering. Notably, the model appears to rely on undesired information. Specifically, it tends to group features based on the number of labels per sample, rather than focusing solely on semantic content. By incorporating LADM for extracting single-class features and employing a predefined ETF classifier to regulate feature distribution across clients, the model learns to focus more on semantic content for clustering. This leads to more disentangled, class-specific feature representations and improved clustering quality. Figure 4 also shows that the pairwise cosine similarity between class-wise average features is substantially reduced, indicating enhanced inter-class separability and stronger discriminative capability. Finally, with the addition of extra regularisation terms during training, the feature clustering becomes more compact and semantically coherent, and the similarity among class-wise average prototypes is further reduced.
Conclusion
@@
This paper investigates multi-label label-skewed FL, a practically important yet underexplored problem setting. This task is particularly challenging due to the complexities inherent in multi-label data, combined with difficulties introduced by the federated system, including heterogeneous local data distributions and inconsistencies between local and global data distributions. Inspired by Neural Collapse theory, we propose a feature disentanglement module and incorporate a predefined ETF structure to facilitate the extraction and clustering of class-specific representations, while also aligning feature distributions across clients in the latent feature space. Furthermore, we incorporate specific regularization losses to suppress noisy negative features and enhance compact class-wise clustering. Experiments on four benchmark datasets across eight settings show that our method outperforms existing approaches, with gains of up to 2.87% in class-wise AUC and 7.29% in class-wise F1 score.
@@ Acknowledgments
Ablation Study (additional)
To complement the results reported in the main manuscript, we present an additional ablation study to evaluate the contribution of each component in our proposed method. The experiments are conducted on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009), with the client data distribution configured using a Dirichlet concentration parameter $\beta=0.5$ and a class-presenting ratio $\gamma=0.5$. The experimental results are summarized in Tables 7 and 8, and Figures 5 and 6. As shown in Table 7, integrating LADM for class-specific feature extraction, along with the use of a predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, leads to performance improvements of 3.38% in class-wise AUC and 3.99% in class-wise F1 score. Additional gains are achieved by introducing regularization terms that enhance the model’s discriminative capacity, resulting in a further class-wise increase of 1.88% in AUC and 4.58% in F1 score. Regarding LADM specifically, the results in Table 8 demonstrate that using fixed, well-designed class-wise queries is generally more effective than learnable queries across most evaluation metrics.
To further analyze model behavior, we visualize the latent feature distributions of the test set using t-SNE under different architectural and training configurations. As shown in Figure 5, incorporating LADM for extracting single-class features, along with a predefined ETF classifier to regulate feature distribution across clients, enables the model to focus on semantic content rather than irrelevant factors such as the number of labels per sample. Furthermore, the addition of regularization terms during training leads to more compact and semantically coherent clusters, while also reducing the similarity among class-wise average prototypes. As shown in Figure 6, the pairwise cosine similarity between class-wise average features decreases with the inclusion of LADM and the ETF classifier, and is further reduced by the added regularization terms, indicating enhanced inter-class separability and stronger discriminative capability.
@@
Table 7: Ablation study of the proposed method on the multi-label CIFAR10 dataset. $\beta$ is set to 0.5 and $\gamma$ is set to 0.5.
@@
Table 8: Ablation study of the class-wise feature extraction block - LADM on the multi-label CIFAR10 dataset.
@@
Figure 5: t-SNE visualisation of test data feature embeddings on the multi-label CIFAR10 experiment with $\beta=0.5$, $\gamma=0.5$. Each colour represents a class. Observing from subfigure (a), without feature disentanglement (LADM), the model appears to rely on undesired information, such as the number of labels per sample, for clustering.
@@
Figure 6: Pair-wise cosine similarity between test data class-wise average feature prototypes on the multi-label CIFAR10 dataset.
$\beta=0.5,\gamma=0.5$
macro-AUC macro-F1 AUC F1
Centralized (Upperbound) $92.20\pm 0.36$ $61.30\pm 0.94$ $90.04\pm 0.65$ $62.02\pm 0.62$
FedAvg (McMahan et al.,, 2017) $82.29\pm 0.46$ $39.47\pm 1.59$ $81.48\pm 0.78$ $40.26\pm 1.09$
FedCurv (Shoham et al.,, 2019) $82.53\pm 0.30$ $39.96\pm 1.28$ $82.10\pm 0.29$ $40.34\pm 1.04$
FedProx (Li et al.,, 2020) $82.45\pm 0.34$ $39.22\pm 0.74$ $81.77\pm 0.48$ $39.66\pm 0.53$
SCAFFOLD (Karimireddy et al.,, 2020) $82.51\pm 0.44$ $39.98\pm 1.41$ $82.08\pm 0.53$ $40.26\pm 1.26$
SphereFed (Dong et al.,, 2022) $\underline{83.63}\pm 1.50$ $42.58\pm 3.20$ $\underline{83.50}\pm 1.87$ $43.18\pm 2.34$
FedLGT (Liu et al.,, 2024) $83.52\pm 0.68$ $\underline{43.60}\pm 1.68$ $83.36\pm 0.71$ $\underline{44.03}\pm 1.71$
[…]
Table 9: Comparisons on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009). Non-IID data distributions across clients are simulated using the Dirichlet concentration parameter ($\beta$) and the class-missing ratio ($\gamma$).
Experiments (additional)
To complement the results reported in the main manuscript and further validate the reproducibility of our method, we present additional experiments on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009). These experiments were repeated three times with different random seeds to evaluate the stability and effectiveness of the proposed approach. The mean and standard deviation of the results are reported in Table 9. As shown in the table, the proposed method demonstrates stable performance across multiple runs and consistently outperforms competing approaches.
Dataset
@@
In this section, we introduce the datasets used in our experiments. To evaluate both the effectiveness and real-world applicability of the proposed method, we conduct experiments on datasets from general computer vision (CV) as well as medical imaging domains. Specifically, we use CIFAR-10 (Krizhevsky et al.,, 2009) and PASCAL VOC (Everingham et al.,, 2010) as general CV datasets, and DermaMNIST (Yang et al., 2023a, ) and ChestX-ray14 (Wang et al.,, 2017) as medical datasets. We simulate the FL system with 10 clients. To model non-IID data distributions, the data is partitioned using a Dirichlet distribution, with the concentration parameter $\beta$ controlling the degree of heterogeneity. To further simulate missing-class scenarios, we constrain the number of classes available to each client using the class-presenting ratio $\gamma$, which specifies the proportion of total classes present locally. Dataset-specific settings are detailed below.
@@
CIFAR-10 (Krizhevsky et al.,, 2009) is a widely used benchmark in CV. It comprises 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. To adapt CIFAR-10 for multi-label classification, following Li et al., 2023a , we create composite samples by combining multiple original images into a single image, with their corresponding labels forming a multi-label ground truth. Specifically, we retain a portion of the original single-label images and augment the dataset by creating an equal number of synthetic samples for each possible pairwise class combination. These composites are constructed by randomly selecting two images from different classes and combining them into one. We evaluate our method on the multi-label CIFAR-10 dataset under two FL configurations, using $\beta=0.5,\ \gamma=0.5$ and $\beta=0.1,\ \gamma=0.5$. Figure 8 illustrates the corresponding class-wise data distribution across clients.
DermaMNIST (Yang et al., 2023a, ) is a skin lesion classification dataset. It consists of 7 diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis-like lesions, dermatofibroma, melanocytic nevi, melanoma, and vascular lesions. To adapt DermaMNIST for multi-label classification, we apply a strategy similar to the one used for multi-label CIFAR-10. Given the long-tailed nature of the original dataset, we preserve its inherent distribution by retaining all single-label samples. Then, we augment the dataset with an equal number of synthetic samples for each possible pairwise label combination. Figure 7 shows examples of the resulting composite multi-label inputs. The resulting dataset comprises the complete set of original samples along with the newly generated multi-label samples. We evaluate our method on this multi-label DermaMNIST dataset under two FL configurations: $\beta=0.5,\gamma=0.71$ and $\beta=0.1,\gamma=0.71$. The corresponding class-wise data distributions across clients are illustrated in Figure 9.
(a) actinic keratoses, vascular lesions
(b) basal cell carcinoma, dermatofibroma
(c) bird, vascular lesions
(d) dermatofibroma, melanoma
@@
PASCAL VOC 2012 (VOC) (Everingham et al.,, 2010) is a widely used benchmark dataset in CV. It consists of approximately 11,500 images, each annotated with one or more object categories selected from a predefined set of 20 classes. These object categories are: aeroplane, bicycle, bird, boat, bottle, bus, car, cat, chair, cow, dining table, dog, horse, motorbike, person, potted plant, sheep, sofa, train, and TV monitor. This dataset poses a challenging multi-label classification problem due to the presence of class imbalance, intra-class variability, and frequent inter-class co-occurrence. For example, the person category often co-occurs with many other object classes. Under a label-skewed FL setting, the non-IID distribution of data and inconsistencies in class co-occurrence patterns across clients further exacerbate the learning difficulty. We evaluate our method on VOC under two FL configurations, using $\beta=0.05$, $\gamma=0.5$ and $\beta=0.01$, $\gamma=0.5$. Figure 10 illustrates the resulting class-wise data distribution across clients.
@@
ChestX-ray14 (Wang et al.,, 2017) is a large-scale medical imaging dataset commonly used for automated thoracic disease classification. It consists of 112,120 frontal-view chest X-ray images collected from 30,805 unique patients. Each image is annotated with one or more disease labels, making the dataset naturally suitable for multi-label classification tasks. These labels were extracted from corresponding radiology reports using natural language processing techniques. The dataset includes 14 disease categories: Atelectasis, Cardiomegaly, Consolidation, Edema, Effusion, Emphysema, Fibrosis, Hernia, Infiltration, Mass, Nodule, Pleural Thickening, Pneumonia, and Pneumothorax. In addition, a No Finding label is used to indicate negative samples in which none of the 14 diseases are present. Due to its scale, real-world variability, and label noise, ChestX-ray14 provides a challenging benchmark for developing and evaluating deep learning models in medical image analysis, particularly in multi-label, imbalanced scenarios. We conduct experiments under two label-skewed FL settings on this dataset, using configurations of $\beta=0.5$, $\gamma=0.5$ and $\beta=0.1$, $\gamma=0.5$. Since a significant portion of the dataset is “No Finding” samples, we distribute these evenly across all clients in both settings. This setup mimics realistic clinical scenarios where healthy cases are common, while disease cases are relatively rare and unequally distributed across institutions. Figure 11 shows the class-wise data distribution across clients.
@@
Figure 8: Distribution of data across local clients used in the experiments on the CIFAR-10 dataset (Krizhevsky et al.,, 2009).
@@
Figure 9: Distribution of data across local clients used in the experiments on the DermaMNIST dataset (Yang et al., 2023a, ).
@@
Figure 10: Distribution of data across local clients used in the experiments on the VOC dataset (Everingham et al.,, 2010).
@@
Figure 11: Distribution of data across local clients used in the experiments on the ChestX-ray14 dataset (Wang et al.,, 2017).

### v3

@@ 5.3 Performance Comparison
ETF Clf LADM $\mathcal{L}_{\text{Neg}}$ $\mathcal{L}_{\text{Pos}}$ macro-AUC macro-F1 micro-AUC micro-F1
@@
query type query init macro-AUC macro-F1 micro-AUC micro-F1
@@
(a) FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + ETF + Dual regularization
Figure 4: Pairwise cosine similarity of class-wise average feature prototypes. Incorporating LADM, ETF-based alignment, and structure-preserving regularization lowers inter-class similarity, reflecting enhanced separability and discrimination.
Figure 5: Examples of attention maps from the multi-label DermaMNIST dataset. Each subfigure shows a sample image alongside the class-specific attention maps generated by the LADM module for the corresponding ground-truth labels. The LADM module captures class-specific features through its attention mechanism, where redder regions indicate areas of higher attention.
5.4 Ablation Study
To evaluate the effectiveness of each component in the proposed method, we conducted an ablation study on the multi-label DermaMNIST dataset. This dataset was selected because it exhibits a strong class imbalance and pronounced inter-class and intra-class variability, which arise from the diverse nature of dermatological conditions. In our experiments, the client data distribution was configured using $\beta=0.1$ and $\gamma=0.71$, which further intensifies the class imbalance and label heterogeneity. The results are summarized in Table 5. As shown in the table, solely using the predefined ETF classifier leads to degraded performance. This is because image-level features often contain multiple semantic cues, and enforcing direct clustering on such entangled representations confuses the model. In contrast, incorporating LADM for class-specific feature extraction, together with the predefined ETF classifier to promote class-wise clustering alignment across clients in the latent space, yields improvements of 0.43% in class-wise AUC and 7.34% in class-wise F1 score. Additional performance gains are achieved by introducing regularisation terms that enhance the model’s discriminative ability, resulting in further improvements of 3.31% in class-wise AUC and 3.43% in class-wise F1 score.
@@ 5.4 Ablation Study
We also investigated the attention mechanism within the LADM module. As shown in Table 6, fixed, well-designed class-wise queries consistently outperform learnable ones across most evaluation metrics. This result is intuitive under label-skewed FL, where clients have distinct local data distributions. When both the queries and the classifier are learnable, models tend to overfit local distributions that deviate from the global one, leading to inconsistencies among clients and degraded global performance after aggregation. In contrast, predefined, well-separated embeddings ensure consistent learning across clients with non-IID data. Figure 5 further illustrates that the LADM attention mechanism effectively focuses on semantically relevant image regions associated with the query class, thereby enabling discriminative class-specific feature extraction.
@@
To further analyse the model’s behaviour, we present t-SNE visualisations of the test data in the latent feature space (Figure 3) and the pairwise cosine similarity between class-wise average features (Figure 4), obtained under different model architectures and training strategies on the multi-label DermaMNIST dataset. As shown in Figure 3(a), when a conventional learnable fully connected (FC) classifier is used, the resulting feature representations exhibit poor clustering. Notably, the model appears to rely on undesired information, grouping features by the number of labels per sample rather than solely by semantic content. By incorporating LADM (Figure 3(b), 4(b)), which extracts single-class features, and further adding a predefined ETF classifier to regulate feature distribution across clients, the model learns to cluster features based on meaningful semantic attributes. This results in improved clustering quality, as evidenced by the substantially reduced pairwise cosine similarity between class-wise average features, indicating enhanced inter-class separability and stronger discriminative capability. Finally, incorporating additional regularization terms during training yields even more compact and semantically coherent feature clusters, with further reductions in prototype similarity (Figure 3(c), 4(c)).
6 Conclusion
@@
This paper addresses the challenging yet underexplored problem of FL under multi-label and label-skewed conditions, where heterogeneous client data lead to severe class imbalance and inconsistent label correlations. To tackle these challenges, we propose FedNCA-ML, a Neural Collapse–inspired framework that enhances feature extraction and alignment across clients with non-IID data distributions. FedNCA-ML integrates a feature disentanglement module with a predefined ETF structure, extending Neural Collapse theory to the multi-label setting and facilitating consistent class-wise representation alignment among clients. Furthermore, two regularization losses are designed to suppress noisy negative features and enhance class-wise compactness in the latent space. Extensive experiments on four datasets under eight different FL settings demonstrate the effectiveness and robustness of the proposed method, yielding consistent improvement of up to 3.92% in class-wise AUC and 4.93% in class-wise F1 score.
@@ Acknowledgments
Appendix A Related Work (Additional)
A.1 Heterogeneous Federated Learning
The foundational Federated Learning (FL) algorithm, FedAvg (mcmahan2017communication), performs iterative weighted averaging of locally trained model parameters, and remains the underlying pipeline for many advanced FL methods. While FedAvg achieves strong performance under independently and identically distributed (IID) data, real-world scenarios often involve non-IID data distributions. Such heterogeneity commonly appears in the form of quantity skew, label-distribution skew, and feature-distribution skew, each of which poses substantial challenges for FL. In the multi-class classification setting, numerous methods have been proposed to address such data heterogeneity within FL settings. FedProx (li2020federated) introduces an L2 regularisation term to constrain local updates and encourage proximity to the global model. FedCurv (shoham2019overcoming) uses Elastic Weight Consolidation (kirkpatrick2017overcoming) to preserve important local parameters to reduce divergence across clients. SCAFFOLD (karimireddy2020scaffold) introduces control variates to correct client drift. SphereFed (dong2022spherefed) projects features onto a hypersphere using a fixed classifier to mitigate discrepancies in latent feature distributions. FedVLS (guo2025exploring) uses global model distillation to alleviate forgetting of missing classes and incorporates loss regularisation to amplify minority class signals, thereby improving performance on underrepresented classes. Although these approaches have demonstrated effectiveness in multi-class settings, their performance tends to deteriorate when directly applied to the more complex multi-label scenario, where each sample may be associated with multiple co-occurring labels. In this work, we focus on addressing the challenges of multi-label FL under quantity-skewed and label-skewed data distributions.
Appendix B Experiment (Additional)
B.1 Ablation Study (Additional)
To complement the results reported in the main manuscript, we present an additional ablation study to evaluate the contribution of each component in our proposed method. The experiments are conducted on the multi-label CIFAR-10 dataset (krizhevsky2009learning), where the client data distribution is configured using a Dirichlet concentration parameter of $\beta=0.5$ and a class-presence ratio of $\gamma=0.5$ (i.e., each client has access to at most 5 out of 10 classes). The experimental results are summarized in Tables 7 and 8, and Figures 6 and 7. As shown in Table 7, integrating LADM for class-specific feature extraction, along with the use of a predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, leads to performance improvements of 3.38% in class-wise AUC and 3.99% in class-wise F1 score. Additional gains are achieved by introducing regularization terms that enhance the model’s discriminative capacity, resulting in a further class-wise increase of 1.88% in class-wise AUC and 4.58% in class-wise F1 score. Regarding LADM specifically, the results in Table 8 demonstrate that using fixed, well-designed class-wise queries is generally more effective than learnable queries across most evaluation metrics.
Table 7: Ablation study of the proposed method on the multi-label CIFAR-10 dataset (krizhevsky2009learning). $\beta$ is set to 0.5 and $\gamma$ is set to 0.5.
@@ B.1 Ablation Study (Additional)
Table 8: Ablation study of the class-wise feature extraction block - LADM on the multi-label CIFAR-10 dataset.
@@
Figure 6: t-SNE visualisation of test data feature embeddings on the multi-label CIFAR-10 experiment with $\beta=0.5$, $\gamma=0.5$. Each colour represents a class. Observing from subfigure (a), without feature disentanglement (LADM), the model appears to rely on undesired information, such as the number of labels per sample, for clustering.
@@
Figure 7: Pair-wise cosine similarity between test data class-wise average feature prototypes on the multi-label CIFAR-10 dataset.
To further analyse model behaviour, we visualize the latent feature distributions of the test set using t-SNE under different architectural and training configurations. As shown in Figure 6, incorporating LADM for feature disentanglement, along with a predefined ETF classifier to regulate feature distribution across clients, enables the model to focus on semantic content rather than irrelevant factors such as the number of labels present in each sample. Furthermore, the addition of regularization terms during training leads to more compact and semantically coherent clusters, while also reducing the similarity among class-wise average prototypes. As illustrated in Figure 7, the pairwise cosine similarity between class-wise average features decreases with the inclusion of LADM and the ETF classifier, and is further reduced by the added regularization terms, indicating enhanced inter-class separability and stronger discriminative capability.
Appendix C Dataset
@@
In this section, we introduce the datasets used in our experiments. To evaluate both the effectiveness and real-world applicability of the proposed method, we conduct experiments on datasets from general computer vision (CV) as well as medical imaging domains. Specifically, we use CIFAR-10 (krizhevsky2009learning) and PASCAL VOC (everingham2010pascal) as general CV datasets, and DermaMNIST (yang2023medmnist) and ChestX-ray14 (wang2017chestxray) as medical datasets. We simulate an FL environment with 10 clients. To model non-IID data distributions, the data is partitioned using a Dirichlet distribution, with the concentration parameter $\beta$ controlling the degree of heterogeneity. To further simulate missing-class scenarios, we constrain the number of classes available to each client using the class presence ratio $\gamma$, which specifies the proportion of total classes present locally. Dataset-specific settings are detailed below.
@@
CIFAR-10 (krizhevsky2009learning) is a widely used benchmark in CV. It comprises 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. To adapt CIFAR-10 for multi-label classification, following li2023neural, we create composite samples by combining multiple original images into a single image, with their corresponding labels forming a multi-label ground truth. Specifically, we retain a portion of the original single-label images and augment the dataset by creating an equal number of synthetic samples for each possible pairwise class combination. Each composite is constructed by randomly selecting two images from different classes and merging them into a single sample. We evaluate our method on the resulting multi-label CIFAR-10 dataset under two FL configurations: $\beta=0.5,\ \gamma=0.5$ and $\beta=0.1,\ \gamma=0.5$. Figure 9 illustrates the corresponding class-wise data distributions across clients.
@@
DermaMNIST (yang2023medmnist) is a skin lesion classification dataset. It consists of 7 diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis-like lesions, dermatofibroma, melanocytic nevi, melanoma, and vascular lesions. To adapt DermaMNIST for multi-label classification, we apply a strategy similar to that used for the multi-label CIFAR-10 dataset. Given the long-tailed nature of the original dataset, we preserve its inherent distribution by retaining all single-label samples. We then augment the dataset by generating an equal number of synthetic samples for each possible pairwise label combination. Examples of the resulting composite multi-label samples are shown in Figure 8. The final dataset comprises the complete set of original samples together with the newly generated multi-label samples. We evaluate our method on this multi-label DermaMNIST dataset under two FL configurations: $\beta=0.5,\gamma=0.71$ and $\beta=0.1,\gamma=0.71$. The corresponding class-wise data distributions across clients are illustrated in Figure 10.
@@
PASCAL VOC (everingham2010pascal) is a widely used benchmark dataset in CV. It contains approximately 11,500 images, each annotated with one or more object categories selected from a predefined set of 20 classes. These categories include: aeroplane, bicycle, bird, boat, bottle, bus, car, cat, chair, cow, dining table, dog, horse, motorbike, person, potted plant, sheep, sofa, train, and TV monitor. This dataset presents a challenging multi-label classification task due to substantial class imbalance, high intra-class variability, and frequent inter-class co-occurrence. For instance, the person class commonly appears alongside many other object categories. Under a label-skewed FL setting, the non-IID data distribution and inconsistent co-occurrence patterns across clients further exacerbate these challenges. We evaluate our method on PASCAL VOC under two FL configurations: $\beta=0.05$, $\gamma=0.5$ and $\beta=0.01$, $\gamma=0.5$. Figure 11 illustrates the resulting class-wise data distributions across clients.
@@
ChestX-ray14 (wang2017chestxray) is a large-scale medical imaging dataset widely used for automated thoracic disease classification. It contains 112,120 frontal-view chest X-ray images collected from 30,805 unique patients. Each image is annotated with one or more disease labels, making the dataset naturally suited to multi-label classification. These labels are extracted from the corresponding radiology reports using natural language processing techniques. The dataset includes 14 disease categories: Atelectasis, Cardiomegaly, Consolidation, Edema, Effusion, Emphysema, Fibrosis, Hernia, Infiltration, Mass, Nodule, Pleural Thickening, Pneumonia, and Pneumothorax. In addition, a No Finding label is used to indicate negative samples in which none of the 14 diseases are present. Due to its scale, real-world variability, and inherent label noise, ChestX-ray14 provides a challenging benchmark for developing and evaluating deep learning models in medical image analysis, particularly in multi-label and imbalanced scenarios. We conduct experiments on this dataset under two label-skewed FL configurations: $\beta=0.5$, $\gamma=0.5$ and $\beta=0.1$, $\gamma=0.5$. Since a significant portion of the dataset (57% of the training data) is No Finding samples, we distribute these samples evenly across all clients in both settings. This setup mimics realistic clinical scenarios where healthy cases are common, while disease cases are relatively rare and unequally distributed across institutions. Figure 12 illustrates the resulting class-wise data distributions across clients.
@@
Figure 9: Distribution of data across local clients in the CIFAR-10 (krizhevsky2009learning) experiments. The class presence ratio ($\gamma$) is set to 0.5 ($\leq$ 5 of 10 classes per client). Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).
@@
Figure 10: Distribution of data across local clients in the DermaMNIST (yang2023medmnist) experiments. The class presence ratio ($\gamma$) is set to 0.71 ($\leq$ 5 of 7 classes per client). Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).
@@
Figure 11: Distribution of data across local clients in the PASCAL VOC (everingham2010pascal) experiments. The class presence ratio ($\gamma$) is set to 0.5 ($\leq$ 10 of 20 classes per client). Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).
@@
Figure 12: Distribution of data across local clients in the ChestX-ray14 (wang2017chestxray) experiments. The ChestX-ray14 dataset contains 14 thoracic disease categories and an additional “No Finding” label. Since a large portion of the dataset (57% of the training data) is “No Finding” samples (i.e., negative cases with all-zero labels), we distribute these samples evenly across all clients to reflect a realistic clinical scenario in which healthy cases are prevalent. The class presence ratio ($\gamma$) is set to 0.5 ($\leq$ 7 of 14 disease classes per client). Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).